OpenStudy (anonymous):

can you help me factoring this? 4(2x-1)² -1

OpenStudy (anonymous):

Difference of squares \(a^{2}-b^{2} = (a-b)(a+b)\) For you, \(a^{2} = 4(2x-1)^{2}\), \(b^{2} = 1\)